the sum of two numbers is 26. three times the second number is two more than the first. what are the two numbers?
Answer by mananth(16946)

the sum of two numbers is 26. three times the second number is two more than the first. what are the two numbers?
..
x+y =26..................1
3y=2+x
3y-x=2...................2
...
add (1)&(2)
4y=28
/4
y =7
the other number = 19
...
m.ananth@hotmail.ca